OpenStudy (anonymous):

solve the equation. x^3 - 1=0. i know it means find the third roots of 1. I have an answer but im not sure if its right. I got 3

OpenStudy (anonymous):

This is powers and roots of complex numbers

OpenStudy (solomonzelman):

\(\large\color{slate}{ a^3-b^3=(a-b)(a^2+ab+b^2) }\)

OpenStudy (solomonzelman):

use that here: \(\large\color{slate}{ x^3-1^3=~..... }\)

OpenStudy (anonymous):

we have to use DeMoivre's Theorem

OpenStudy (solomonzelman):

I am out.... (why ?)

OpenStudy (anonymous):

idk, thats how we do it

OpenStudy (anonymous):

for ex: [32(cos(2pi) + isin(2pi)]^1/5